OpenStudy (imterribleatmath):

given f(x)=x^2+7x and g(x)=x^4 choose the expression for (f g) (x) @3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (will.h):

You mean f(g(x)?

OpenStudy (3mar):

What do you think?

OpenStudy (imterribleatmath):

like i said in just as clueless on this as the last

OpenStudy (will.h):

Well no you didn't say anything

OpenStudy (3mar):

@Will.H It is your turn as you want to answer!

OpenStudy (imterribleatmath):

@Will.H

OpenStudy (will.h):

Can't help you as I don't know what (F g) (x) means

OpenStudy (will.h):

I'll tag a spectacular helper @563blackghost

OpenStudy (3mar):

Can I proceed?

OpenStudy (imterribleatmath):

is there a way to post a picture ?

OpenStudy (will.h):

Yeah. Take a screenshots prnt.SC

OpenStudy (imterribleatmath):

@Will.H

OpenStudy (will.h):

Alright so here's what I believe. It's a multiplication process they want to multiply function f and g (x^4)(x^2+ 7x) Now use destrubtive property When multiplying here the exponents be added X^6 + 7x^5 but since that doesn't match any of your options I'd assume This would bw the process (x^2 + 7x)^4 and the didn't want you to simplify so it would be C

OpenStudy (will.h):

Again am not certain.

OpenStudy (imterribleatmath):

your saying it's (x^2+7x)^4

OpenStudy (imterribleatmath):

just to be sure.

OpenStudy (will.h):

Probably. And let me know if it's correct once you submit it

OpenStudy (imterribleatmath):

it was incorrect. it's ok

OpenStudy (will.h):

Sorry about that. What was the correct option

OpenStudy (imterribleatmath):

it doesn't show

OpenStudy (will.h):

Okay.

OpenStudy (imterribleatmath):

this is the next one if u wanna take a shot

OpenStudy (will.h):

No I hate answering things based on speculations. Probably someone else will help if you open a new thread

OpenStudy (imterribleatmath):

ok thanks anyways

OpenStudy (will.h):

Tag zepdrix when u need help

OpenStudy (osprey):

f(x)=x^2+7x and g(x)=x^4 choose the expression for (f g) (x) (fg)(x)=(x^2+7x).(x^4) I think.